Marginal tax rates Partner A, a single taxpayer, is one of two partners in a small business. As such, she receives pass-through income that is taxed at her personal tax rates. After all adjustments and deductions have been made, including the 20% qualified business income deduction, she is preparing o calculate her axes owed or the year. Using the tax rate schedule grven here perform he ollowing: a. Find the marginal tax rate for the following levels of sole proprietorship earnings before taxes: $13,600; S60,000: $89,900; $150,000; $248,000; $454,000; and $1.4 million. b. Plot the marginal tax rates (measured on the y-axis) against the pretax income levels (measured on the x-axis). Explain the relationship between these variables (Round to the nearest integer.) a. The marginal tax rate for eamings before taxes of $13,600 is The total taxes due for earnings before taxes of $13,600 is $ 1, (Round to the nearest dollar.) The marginal tax rate for eamings before taxes of $60,000 is Round to the nearest integer.) The total taxes due for earnings before taxes of $60,000 is (Rou More Info The marginal tax rate for earnings before taxes of $89,900 is The total taxes due for earnings before taxes of $89,900 is $. (Rou Tax Rate - Single Taxpayers -2018 Tax calculation The marginal tax rate for earnings before taxes of $150,000 is Range of taxable income (Marginal rate x amount over base bracket) Base tax S0 to S9,525 $0 (10% (12% (22% (24% (32% (35% (37% S0) 9,525) 38,700) 82,500) 157,500) 200,000) 500,000) x amount over + The total taxes due for earnings before taxes of $150,000 is$(Ro 9,525 to 38,700 953 x amount over + 38,700 to 82,500 4.454 x amount over + The marginal tax rate for earnings before taxes of $248,000 is 82,500 to 157,500 14,090 x amount over + 157,500 to 200,000 32,090 x amount over + The total taxes due for earnings before taxes of $248,000 is$(Ro 200,000 to 500,000 45,690 150,690 x amount over + Over 500,000 x amount over + The marginal tax rate for earnings before taxes of $454,000 is . ( The total taxes due for earnings before taxes of $454,000 is (R Print Done The marginal tax rate for earnings before taxes of $1.4 million is 1 %. The total taxes due for earnings before taxes of $1.4 million is SD (Round to the nearest dollar.) Click to select your answerfs)
The correct shape of the curve is: (Select the best answer belaw.) ов. 25 20- 20- 15 10 O 200 400 600 800 1,0001.2001.400 500 1,000 Pretax Income ($000) Pretax Income ($000) с. O D. 20 Click to select your answer(s)
Explain the relationship between these variables. (Select from the drop-down menus.) As income increases to ▼1, the marginal tax rate approaches and peaks at
Pretax Income (000) Pretax Income ($000) O c. O D. 30- 10 500 1,000 Pretax Income ($000) 1,000 0 Pretax Income ($000)
Answer #1

Part A

Marginal tax rate for earnings before taxes of $13600 = 12%

Total taxes due for earnings before taxes of $13600 = 953+(12%*(13600-9525)) = $1442

Marginal tax rate for earnings before taxes of $60000 = 22%

Total taxes due for earnings before taxes of $60000 = 4454+(22%*(60000-38700)) = $9140

Marginal tax rate for earnings before taxes of $89900 = 24%

Total taxes due for earnings before taxes of $89900 = 14090+(24%*(89900-82500)) = $15866

Marginal tax rate for earnings before taxes of $150000 = 24%

Total taxes due for earnings before taxes of $150000 =14090+(24%*(150000-82500)) = $                30290

Marginal tax rate for earnings before taxes of $248000 = 35%

Total taxes due for earnings before taxes of $248000 = 45690+(35%*(248000-200000)) = $62490

Marginal tax rate for earnings before taxes of $454000 = 35%

Total taxes due for earnings before taxes of $454000 = 45690+(35%*(454000-200000)) = $134590

Marginal tax rate for earnings before taxes of $1.4 million = 37%

Total taxes due for earnings before taxes of $1.4 million = 150690+(37%*(14000000-500000)) = $5145690

Part B

Option B

Chart B represents the following sequence of marginal taxes

Part C

As income increases to $500,000 the marginal tax rate approaches and peaks at 37%